OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y=2/x^2 +9

OpenStudy (anonymous):

you have lots of options for this one could be \(g(x)=x^2+9\) and \(f(x)=\frac{2}{x}\)

OpenStudy (anonymous):

I need an explanation to prove how they equal x. I know I need to plug them in, I just don't know exactly how @satellite73

OpenStudy (anonymous):

what does "how they equal \(x\) " mean? \(x\) is the variable

OpenStudy (anonymous):

Sorry, I was thinking of a different problem im working on xD

OpenStudy (anonymous):

if \(g(x)=x^2+9\) and \(f(x)=\frac{2}{x}\) then \[f(g(x))=f(x^2+9)=\frac{2}{x^2+9}\]